# Canonicalization for Unreproducible Builds in Java

### v1

@@ Abstract.
The increasing complexity of software supply chains and the rise of supply chain attacks have elevated concerns around software integrity. Users and stakeholders face significant challenges in validating that a given software artifact corresponds to its declared source. Reproducible Builds address this challenge by ensuring that independently performed builds from identical source code produce identical binaries. However, achieving reproducibility at scale remains difficult, especially in Java, due to a range of non-deterministic factors and caveats in the build process. In this work, we focus on reproducibility in Java-based software, archetypal of enterprise applications. We introduce a conceptual framework for reproducible builds, we analyze a large dataset from Reproducible Central, and we develop a novel taxonomy of six root causes of unreproducibility. We study actionable mitigations: artifact and bytecode canonicalization using OSS-Rebuild and jNorm respectively. Finally, we present Chains-Rebuild, a tool that raises reproducibility success from 9.48% to 26.89% on 12,283 unreproducible artifacts. To sum up, our contributions are the first large-scale taxonomy of build unreproducibility causes in Java, a publicly available dataset of unreproducible builds, and Chains-Rebuild, a canonicalization tool for mitigating unreproducible builds in Java.
@@ 1. Introduction
The growing complexity of software supply chains (Soto-Valero et al., 2022; Cox, 2019), coupled with the increasing frequency of supply chain attacks  [^1] https://www.sonatype.com/hubfs/SSCR-2024/SSCR_2024-FINAL-10-10-24.pdf, raises concerns about software integrity. In such a fragmented ecosystem, relying solely on assumptions about the identity of the distributor (Thompson, 1984) is no longer sufficient - what is needed is verifiable evidence that the software one installs corresponds exactly to its declared source. This challenge is especially acute in open source environments, where binaries are often distributed separately from their source code (Davies et al., 2011), making it difficult for users to independently validate what they are running. As a result, the focus is shifting toward techniques that can guarantee that what is built matches what was intended, regardless of who performs the build.
@@
This technique is formally called “Reproducible Build” (Williams et al., 2025; Cox, 2025): builds are considered “reproducible” if and only if the build process is deterministic, where the same binary can be computed from the same source code by independent parties (Lamb and Zacchiroli, 2022). It helps prevent attacks on the build process, where an attacker modifies it to insert backdoors or malicious code into the software application to be built (Perry, 2013). Any backdoor or malicious code would be detected by the verifier as the built artifact would not match the original artifact. This approach has already seen adoption in security-critical environments, such as in certain federal government contracts, where vendors are required to supply source code so the agencies can perform their own builds and verify the integrity of the software  [^2] https://news.ycombinator.com/item?id=43492115.
Despite the promises of Reproducible Builds, ensuring and verifying reproducibility at scale remains technically challenging due to the multitude of spurious differences in build outputs (e.g., timestamps, file ordering). Those differences make binaries appear different even when built from the same source, with an untampered build pipeline (Fourné et al., 2023). Spurious differences hinder reproducibility and make it difficult to justify the differences between two builds. XZ Utils is a widely used compression library that is used in almost all Linux distributions. In 2024, it is found that the XZ Utils has a backdoor to give remote code execution to the attacker (Przymus and Durieux, 2025). However, this backdoor is only observed in the official tarballs on the registry and do not exist in the git repository. Such supply chains attacks can be prevented if build process is reproducible. In the case of XZ Utils, if the tarball on the package registry does not match the one built from the source code, then the user can be sure that the tarball is tampered with (Malka, 2025).
@@
In this work, we contribute to solving the problem of achieving build reproducibility in the context of enterprise software in Java. Recall that Java has consistently been one of the top 5 programming languages in the world for the past 5 years  [^3] https://innovationgraph.github.com/global-metrics/programming-languages, underscoring its widespread use and relevance. Its importance in the context of finance (Soto-Valero et al., 2022; Morales, 2021), government services (Robinson, 2017), and military applications  [^4] http://pdf.cloud.opensystemsmedia.com/vita-technologies.com/Aonix.Jun07.pdf  [^5] https://tsri.com/news-blog/press/u-s-department-of-defense-mainframe-cobol-to-java-on-aws highlights the need for reproducible builds in Java. We aim at building the most comprehensive taxonomy of unreproducible builds in Java, and the corresponding mitigation strategies. Furthermore, we perform a deep study how canonicalization - removal of non-deterministic differences - is a good solution for mitigating unreproducibility.
@@
Our study reveals 6 original categories of unreproducibility causes in Java, each with the underlying root causes and the best practice mitigations. For example, changes in the bytecode in Java artifacts. Our large-scale experiments empirically measure the effectiveness of two state-of-the-art canonicalization techniques. First, we show that bytecode canonicalization with jNorm (Schott et al., 2024) can successfully canonicalize 267 out of 898 unreproducible artifacts with JVM bytecode. We study the reasons why jNorm fails to canonicalize the remaining artifacts, which are promising areas of future work. Next, we evaluate artifact canonicalization using Google’s OSS-Rebuild and find that it can successfully canonicalize 9.48% (1,165 / 12,283) of unreproducible artifacts. By implementing new canonicalization patterns informed by our taxonomy, our new tool Chains-Rebuild increases the verification success rate to 26.89% (3,303 / 12,283). Our paper is the first to demonstrate, at a large scale, that canonicalization is a promising mitigation strategy for fixing build unreproducibility in the Java ecosystem.
@@
  • A large-scale experiment on the effectiveness of canonicalization, showing that it is a promising solution to mitigate unreproducibility in Java. Artifact canonicalization is more effective at scale than bytecode canonicalization.
  • A publicly available, curated dataset of 7,961 unreproducible Maven releases for future research.
@@ 2.1. Definitions
Developers often release reusable software in package registries, making them widely available for integration into third-party software projects. For our work on build reproducibility, we need to introduce key terminology. We work in the context of the Java programming language and its Maven build system.
@@
Module: A module is a uniquely addressable component in a software system. In Maven, a module is referenced by a combination of Group Identifier, Artifact Identifier, and Version. Group ID is a unique identifier for the organization, eg. “org.apache”   [^6] https://maven.apache.org/guides/mini/guide-reproducible-builds.html. Artifact ID is the unique name of the module and is unique within the group ID. For example, com.google.guava:guava:32.0.0-jre is a Maven module where com.google.guava is the group ID, guava is the artifact ID, and 32.0.0-jre is the version. There are two types of modules: a source module is a module that is hosted on a version control system (eg. Github), not necessarily published on a package registry; a released module is one published on a package registry, Maven Central in this paper. Both released and source modules can be identified by the combination of group ID, artifact ID, and version to refer to both types of module. In the latter case, we assume the presence of a mapping from version to Git tag or Git commit.
Project: A Maven project is a collection of source modules in the same source repository. It contains build configuration for all the modules in a so-called ‘POM file’. For example, Maven project guava contains 6 source modules - guava-parent, guava-tests, guava-testlib, guava-gwt, guava-bom, and guava.
@@ 2.2. Build Reproducibility
(1) $ReferenceArtifact\leftarrow Build(SourceModule)$
(2) $RebuildArtifact\leftarrow Rebuild(SourceModule)$
(3) $ReferenceArtifact=RebuildArtifact$
There are 2 other terms in the literature that are used interchangeably with build reproducibility - “verifiable builds” and “accountable builds” (Pohl et al., 2025). Builds are verifiable if the build process can be modified in order to make the output artifact bit-by-bit identical. Accountable builds are the same as verifiable builds, but instead of removing the differences, these differences are explained and documented so that unreproducibility can be understood upon inspection. In our paper, we stick with the term “build reproducibility” as the other two terms are extensions of build reproducibility with extra steps.
@@ 2.3. Mitigation
We claim that any cause of unreproducibility can be mitigated in three ways - fixing the build process, fixing the rebuild process, or canonicalizing the output of the rebuild process.
@@ 3.1. The Reproducible Central Project
The Reproducible Central project (Boutemy, 2024) is a rebuild infrastructure to verify whether Maven projects are reproducible. It specifies list of steps to build a Java project from source in a buildspec file  [^7] https://github.com/jvm-repo-rebuild/reproducible-central/blob/master/doc/BUILDSPEC.md. Then, the Reproducible Central script takes this buildspec as input to rebuild a project. It produces buildinfo and buildcompare files; buildinfo records the output of the build in terms of the artifact names, sizes, and checksums (Builds, 2024); buildcompare reports which artifacts are reproducible and which are unreproducible.
We use the Reproducible Central project dataset as a starting point to study the causes of unreproducible builds in Java. We take a snapshot of the Reproducible Central dataset on 8th October, 2024 whose commit is d280bf1. At this date, Reproducible Central contains 706 Maven projects, their corresponding build scripts, and a history of 4,956 releases. It is the largest dataset attempting rebuilds of Java projects and is also maintained by maintainers of Apache Maven.
Our goal is to curate a dataset of unreproducible released Maven modules in order to study the reasons why unreproducible builds occur. In order to curate the dataset of unreproducible released modules, we first apply 2 filters before running the rebuilds and 1 filter afterwards.
@@ 3.2. Filtering
3.2.1. Maven Projects
@@
First, we filter out Maven projects that require manual intervention to build so that we can run the entire rebuild process automatically. This is indicated in buildspec file either by the presence of build tool or when the rebuild command is prefixed with ‘SHELL’. The rebuild script would spawn an interactive terminal when any of the above conditions are met. We exclude projects that do not have a POM file at the root directory as it also requires manual intervention to get the path to rebuild artifacts in build directory. Second, we only include projects that use Maven as their build tool (some projects use Gradle or SBT). This is because Maven is the most mature build tool for Java projects, most widely used in the Java ecosystem (Sulír and Porubän, 2016). Moreover, Maven builds success rate is higher than Gradle and Ant (Maes-Bermejo et al., 2022). Once all rebuilds are done, we exclude projects where the unreproducible artifacts are not a subset of the all artifacts released on Maven Central. This can happen when not all artifacts produced by the build process are released on Maven Central. For example, content/org.infinispan:protostream-aggregator:5.0.7.Final produces an artifact rotostream-integrationtests-5.0.7.Final-test-sources.jar which is not released on Maven Central. After applying the filters, we end up with 3,927 releases over 665 Maven projects.
3.2.2. Released Modules
@@
Next, we need to identify the released modules to compare against. In order to map each unreproducible project to corresponding released modules, we first identify the Maven modules in each project. org.apache.aries.cdi:org.apache.aries.cdi.executable:1.1.5 is an example of a source module that does not have a corresponding URL on Maven Central. We parse the POM file located at the root directory of the project as it contains the required information about the modules. For example, jpmml-python is a Maven project hosted on GitHub  [^8] https://github.com/jpmml/jpmml-python/tree/1.2.2 which has 3 Maven modules with version 1.2.2 - org.jpmml:jpmml-python, org.jpmml:jpmml-python:1.2.2, and org.jpmml:pmml-python. Thus, jpmml-python Maven project on Reproducible Central corresponds to 3 Maven modules in our dataset. 2 out of 3 released modules in this project contain unreproducible artifacts. org.jpmml:pmml-python-testing:1.2.2 has pmml-python-testing-1.2.2-sources.jar and org.jpmml:pmml-python:1.2.2 has pmml-python-1.2.2-sources.jar. Finally, there are 4 cases where unreproducible reference and rebuild artifact are not mapped correctly. Maven module org.apache.ratis:ratis-assembly:3.1.0 and 3 more of its versions have an artifact with .pom extension. The rebuild script maps this artifact to the Jar file produced during rebuild. This is a bug in the Reproducible Central rebuild script and have reported this to the author  [^9] https://github.com/jvm-repo-rebuild/reproducible-central/issues/2164.
@@
At the end of the pipeline, we end up with 7,961 Maven modules which have a total of 34,271 artifacts. Out of 34,271 artifacts, there are 12,283 unreproducible artifacts (35.8%).
@@ 3.3. Description of the Dataset
The dataset contains groupId, artifactId, version, and the URL of all 7,961 unreproducible released modules on Maven Central. It also contains the reference and rebuild artifact pair of all 12,283 unreproducible artifacts. It is available at https://github.com/chains-project/reproducible-central/. To our knowledge, this is the largest ever curated dataset of unreproducible artifacts in the context of Java. Based on the dataset, we can answer the following research questions:

### v2

# Causes and Canonicalization of Unreproducible Builds in Java

@@ Abstract.
The increasing complexity of software supply chains and the rise of supply chain attacks have elevated concerns around software integrity. Users and stakeholders face significant challenges in validating that a given software artifact corresponds to its declared source. Reproducible Builds address this challenge by ensuring that independently performed builds from identical source code produce identical binaries. However, achieving reproducibility at scale remains difficult, especially in Java, due to a range of non-deterministic factors and caveats in the build process. In this work, we focus on reproducibility in Java-based software, archetypal of enterprise applications. We introduce a conceptual framework for reproducible builds, we analyze a large dataset from Reproducible Central, and we develop a novel taxonomy of six root causes of unreproducibility. We study actionable mitigations: artifact and bytecode canonicalization using OSS-Rebuild and jNorm respectively. Finally, we present Chains-Rebuild (improvements to OSS-Rebuild), a tool that raises reproducibility success from 9.48% to 26.60% on 12,803 unreproducible artifacts. To sum up, our contributions are the first large-scale taxonomy of build unreproducibility causes in Java, a publicly available dataset of unreproducible builds, and Chains-Rebuild, a canonicalization tool for mitigating unreproducible builds in Java.
@@ 1. Introduction
The growing complexity of software supply chains (Soto-Valero et al., 2022; Cox, 2019), coupled with the increasing frequency of supply chain attacks  [^1] https://www.sonatype.com/hubfs/SSCR-2024/SSCR_2024-FINAL-10-10-24.pdf, raises concerns about software integrity. In such a fragmented ecosystem, relying solely on assumptions about the identity of the distributor (Thompson, 1984) is no longer sufficient — what is needed is verifiable evidence that the software one installs corresponds exactly to its declared source. This challenge is especially acute in open source environments, where binaries are often distributed separately from their source code (Davies et al., 2011), making it difficult for users to independently validate what they are running. As a result, the focus is shifting toward techniques that can guarantee that what is built matches what was intended, regardless of who performs the build.
@@
This technique is formally called “Reproducible Build” (Williams et al., 2025; Cox, 2025): builds are considered “reproducible” if and only if the build process is deterministic, where the same binary can be computed from the same source code by independent parties (Lamb and Zacchiroli, 2022). It helps prevent attacks on the build process, where an attacker modifies it to insert backdoors or malicious code into the software application to be built (Perry, 2013). Any backdoor or malicious code would be detected by the verifier as the built artifact would not match the original artifact. This approach has already seen adoption in security and privacy critical environments, such as Bitcoin and Tor Browser (Levsen et al., 2025). Ensuring that clients for both of them are free of any backdoors injected by the build system.
Despite the promises of Reproducible Builds, ensuring and verifying reproducibility at scale remains technically challenging due to the multitude of spurious differences in build outputs (e.g., timestamps, file ordering). Those differences make binaries appear different even when built from the same source, with an untampered build pipeline (Fourné et al., 2023). Spurious differences hinder reproducibility and create a new attack surface. Several high-profile supply chain attacks have exploited unreproducible builds. For instance, in 2024 it was discovered that the official tarballs of XZ Utils, compression library used in almost all Linux distributions, contained a backdoor that enabled remote code execution, even though the backdoor was absent in the project’s Git repository (Przymus and Durieux, 2025). Similarly, the Ultralytics Python package, a widely used library for AI models, was compromised when attackers injected malicious code into the build pipeline, resulting in a PyPI release with a backdoor (Larson, 2024). A related incident targeted the Solana ecosystem as well (Claburn, 2024). These incidents share a common vector: the artifacts distributed through registries or release tarballs differed from what could be obtained by building directly from the publicly available source code. If the build processes for these projects had been reproducible, users or third party rebuilders could have verified that the distributed binaries and packages matched the source. Any discrepancies would have immediately signaled tampering, thereby preventing such attacks (Malka, 2025).
@@
In this work, we contribute to solving the problem of achieving build reproducibility in the context of enterprise software in Java. Recall that Java has consistently been one of the top 5 programming languages in the world for the past 5 years  [^2] https://innovationgraph.github.com/global-metrics/programming-languages, underscoring its widespread use and relevance. Its importance in the context of finance (Soto-Valero et al., 2022; Morales, 2021), government services (Robinson, 2017), and military applications  [^3] http://pdf.cloud.opensystemsmedia.com/vita-technologies.com/Aonix.Jun07.pdf  [^4] https://tsri.com/news-blog/press/u-s-department-of-defense-mainframe-cobol-to-java-on-aws highlights the need for reproducible builds in Java. We aim at building the most comprehensive taxonomy of unreproducible builds in Java, and the corresponding mitigation strategies. Furthermore, we perform a deep study how canonicalization — removal of non-deterministic differences — is a good solution for mitigating unreproducibility.
@@
Our study reveals 6 reasons for unreproducibility in Java, each with the underlying root causes and the best practice mitigations. For example, changes in the bytecode in Java artifacts. Our large-scale experiments empirically measure the effectiveness of two state-of-the-art canonicalization techniques. First, we show that bytecode canonicalization with jNorm (Schott et al., 2024) can successfully canonicalize 275 out of 936 (29.38%) unreproducible artifacts with JVM bytecode. We study the reasons why jNorm fails to canonicalize the remaining artifacts, which are promising areas of future work. Next, we evaluate artifact canonicalization using Google’s OSS-Rebuild and find that it can successfully canonicalize 9.41% (1,205 / 12,803) of unreproducible artifacts. By implementing new canonicalization patterns informed by our taxonomy, our new tool Chains-Rebuild increases the verification success rate to 26.60% (3,405/ 12,803). Our paper is the first to demonstrate, at a large scale, that canonicalization is a promising mitigation strategy for fixing build unreproducibility in the Java ecosystem.
@@
  • A large-scale experiment on the effectiveness of canonicalization, showing that it is a promising solution to mitigate unreproducibility in Java, with artifact canonicalization being more effective at scale than bytecode canonicalization.
  • A publicly available, curated dataset of 8,292 unreproducible Maven releases for future research.
@@ 2.1. Definitions
Developers often release reusable software in package registries, making them widely available for integration into third-party software projects. For our work on build reproducibility, we need to introduce key terminology. We work in the context of the Java programming language and its Maven build system. In the following, we illustrate the key concepts with the example of Apache Guava.
@@
Module: A module is a uniquely addressable component in a software system. In Maven, a module is referenced by a combination of Group Identifier, Artifact Identifier, and Version. Group ID is a unique identifier for the organization, e.g., “org.apache”   [^5] https://maven.apache.org/guides/mini/guide-reproducible-builds.html. Artifact ID is the unique name of the module and is unique within the group ID. For example, com.google.guava:guava:32.0.0-jre is a Maven module where com.google.guava is the group ID, guava is the artifact ID, and 32.0.0-jre is the version. There are two types of modules: a source module is a module that is hosted on a version control system (e.g., Github), not necessarily published on a package registry; a released module is one published on a package registry, Maven Central in this paper. Both released and source modules can be identified by the combination of group ID, artifact ID, and version. In the latter case, we assume the presence of a mapping from version to Git tag or Git commit.
Project: A Maven project is a collection of source modules in the same source repository. It contains build configuration for all the modules in a so-called ‘POM file’. For example, the Maven project guava contains 6 source modules — guava-parent, guava-tests, guava-testlib, guava-gwt, guava-bom, and guava.
@@ 2.2. Build Reproducibility
[…]
There are two other terms in the literature that are used interchangeably with build reproducibility — “verifiable builds” and “accountable builds” (Pohl et al., 2025). Builds are verifiable if the build process can be modified in order to make the output artifact bit-by-bit identical. Accountable builds are the same as verifiable builds, but instead of removing the differences, these differences are explained and documented so that unreproducibility can be understood upon inspection. In our paper, we stick with the term “build reproducibility” as the other two terms are extensions of build reproducibility with extra steps.
@@ 2.3. Mitigation
Our background study on fixing reproducibility (see  Subsection 8.3), shows that any cause of unreproducibility can be mitigated in three ways — fixing the build process, fixing the rebuild process, or canonicalizing the output of the rebuild process.
@@ 3. Dataset of Unreproducible Builds
Dataset Stage Projects Releases
Count Count
Initial Reproducible Central (d280bf1) 706 4,956
After filters in Subsubsection 3.2.1 684 4,030
After filters in Subsubsection 3.2.2 271 1,144
Table 1. Evolution of dataset size through filtering stages, showing the reduction from initial Reproducible Central data to final unreproducible releases.
@@ 3.1. The Reproducible Central Project
The Reproducible Central project (Boutemy, 2024) is a rebuild infrastructure to verify whether Java projects are reproducible. It specifies list of steps to build a Java project from source in a buildspec file  [^6] https://github.com/jvm-repo-rebuild/reproducible-central/blob/master/doc/BUILDSPEC.md. Then, the Reproducible Central script takes this buildspec as input to rebuild a project. It produces buildinfo and buildcompare files; buildinfo records the output of the build in terms of the artifact names, sizes, and checksums (Builds, 2024); buildcompare reports which artifacts are reproducible and which are unreproducible.
We use the Reproducible Central project dataset as a starting point to study the causes of unreproducible builds in Java. We take a snapshot of the Reproducible Central dataset on 8th October, 2024 whose commit is d280bf1. At this date, Reproducible Central contains 706 projects, their corresponding build scripts, and a history of 4,956 releases. It is the largest dataset attempting rebuilds of Java projects and is also maintained by maintainers of Apache Maven.
Our goal is to curate a dataset of unreproducible released Java modules in order to study the reasons why unreproducible builds occur. In order to curate the dataset of unreproducible released modules, we first apply one filter before running the rebuilds and two filter afterwards.
@@ 3.2. Filtering
Table 1 shows the evolution of the dataset size through filtering stages. We start with 706 Java projects and 4,956 releases in the Reproducible Central dataset. It contains 4,851, 103, and 2 releases built with Maven [^7] https://maven.apache.org/, Gradle [^8] https://gradle.org/, and SBT  [^9] https://www.scala-sbt.org/, respectively. Next, we apply two filters to the dataset which reduces the dataset to 684 Java projects and 4,030 releases. Finally, we only keep the unreproducible releases and projects which results in 271 Java projects and 1,144 releases.
Next, we account for the fact that the 271 Java projects are made of multiple modules. Consequently, we further split the 271 Java projects and their 1,144 releases into 8,292 unreproducible module releases as we account for unreproducibility per <GroupId, ArtifactId, Version> coordinates of Java modules rather than the entire source project.
3.2.1. Filter Java Projects
@@
We filter out Java projects that require manual intervention to build. This is to ensure that we can run the entire rebuild process automatically. This is indicated in buildspec file either by the presence of build tool or when the rebuild command is prefixed with ‘SHELL’. The rebuild script would spawn an interactive terminal when any of the above conditions are met. We exclude projects that do not have a POM file at the root directory as it also requires manual intervention to get the path to rebuild artifacts in build directory. Once all rebuilds are done, we apply the second filter that excludes projects where the unreproducible artifacts are not a subset of all the artifacts released on Maven Central. This can happen when not all artifacts produced by the build process are released on Maven Central. For example, content/org.infinispan:protostream-aggregator:5.0.7.Final produces an artifact rotostream-integrationtests-5.0.7.Final-test-sources.jar which is not released on Maven Central. After applying the two filters, we end up with 4,030 releases over 684 Java projects.
3.2.2. Filter Released Modules
@@
Next, we need to identify the released modules to compare against. In order to map each unreproducible project to corresponding released modules, we first identify the Java submodules in each project. org.apache.aries.cdi:org.apache.aries.cdi.executable:1.1.5 is an example of a source module that does not have a corresponding URL on Maven Central. We parse the POM file located at the root directory of the project as it contains the required information about the modules. For example, jpmml-python is a Maven project hosted on GitHub  [^10] https://github.com/jpmml/jpmml-python/tree/1.2.2 which has three Maven modules with version 1.2.2org.jpmml:jpmml-python, org.jpmml:pmml-python:1.2.2, and org.jpmml:pmml-python-testing. Thus, jpmml-python Maven project on Reproducible Central corresponds to three Maven modules in our dataset. Two out of three released modules in this project contain unreproducible artifacts. org.jpmml:pmml-python-testing:1.2.2 has pmml-python-testing-1.2.2-sources.jar and org.jpmml:pmml-python:1.2.2 has pmml-python-1.2.2-sources.jar. Finally, there are four 4 cases where unreproducible reference and rebuild artifact are not mapped correctly. Maven module org.apache.ratis:ratis-assembly:3.1.0 and three more of its versions have an artifact with .pom extension. The rebuild script maps this artifact to the Jar file produced during rebuild. The third and final filter excludes these four cases. This is a bug in the Reproducible Central rebuild script and we have reported this to the author  [^11] https://github.com/jvm-repo-rebuild/reproducible-central/issues/2164.
@@
At the end of the pipeline, we end up with 8,292 released Java modules which have a total of 35,956 artifacts from 4,030 releases of 684 projects. Finally, we end up with 271 Java projects and 1,144 releases which are unreproducible and they correspond to 8,292 unreproducible released modules that have 12,803/35,956 (35.8%) unreproducible artifacts.
@@ 3.3. Description of the Dataset
The dataset contains groupId, artifactId, version, and the URL of all 8,292 unreproducible released modules on Maven Central. It also contains the reference and rebuild artifact pair of all 12,803 unreproducible artifacts. It is available at https://github.com/chains-project/reproducible-central/. To our knowledge, this is the largest ever curated dataset of unreproducible artifacts in the context of Java. Based on the dataset, we can answer the following research questions: